OpenStudy (anonymous):

Find the total area for the cylinder with the given measurement. r = 2, h = 4

OpenStudy (anonymous):

Volume or surface area? Remember, the former is the product of the base area and height, whereas the latter is two times the area of the circular surface, times the product of the circumference and height.

OpenStudy (paxpolaris):

total surface area\[\Large=2\pi r \left( r+h \right)=24\pi\]